OpenStudy (lgbasallote):

Find the area of the circle that has a center at point (2,3) and tangent to the line 4x - 3y = 5

myininaya (myininaya):

Ok so lets put down what we know: \[(x-2)^2+(y-3)^2=r^2\] where r is the radius this is the thing we want to find if we can find r then we can find the area \[A=\pi r^2\] We also know that the line 4x-3y=5 is tangent to the circle So at some point, (a,b) , on the circle is the line tangent to the circle Solving 4x-3y=5 for y we get \[-3y=-4x+5\] \[y=\frac{4}{3}x-\frac{5}{3}\] Where 4/3 is the slope of this line So when is y'=4/3 \[2(x-2)+2(y-3)y'=0\] \[x-2+(y-3)y'=0 \text{ note: I divided both sides by 2}\] \[(y-3)y'=-x+2\] \[y'=\frac{-x+2}{y-3}\] \[\frac{4}{3}=\frac{-a+2}{b-3} \text{ but gosh darn it we have two unknowns in this equation}\] Remember at that point (a,b) did we have the slope is 4/3 Need more thinking ...

myininaya (myininaya):

we also know that \[(a-2)^2+(b-3)^2=r^2 \] and \[b=\frac{4}{3}a-\frac{5}{3}\]

OpenStudy (anonymous):

pi*(1.2)^2

OpenStudy (lgbasallote):

lol wow derivatives in an analytic geometry question :P im *kind of* following though :D

OpenStudy (anonymous):

derivatives?!!!

myininaya (myininaya):

So \[\frac{4}{3}=\frac{-a+2}{\frac{4}{3}a-\frac{5}{3}-3} \cdot \frac{3}{3}\] \[\frac{4}{3}=\frac{3(-a+2)}{4a-5-9}\] \[\frac{4}{3}=\frac{3(-a+2)}{4a-14}\] \[4(4a-14)=3(-3a+6)\] \[16a-56=-9a+18\] \[25a=74\] \[a=\frac{74}{25}\] \[b=\frac{4}{3}a-\frac{5}{3}=\frac{4}{3}(\frac{74}{25})-\frac{5}{3}\] \[b=\frac{296}{75}-\frac{125}{75}=\frac{171}{75}\] So we have \[(a-2)^2+(b-3)^2=r^2\] \[(\frac{74}{25}-2)^2+(\frac{171}{75}-3)^2=r^2\]

myininaya (myininaya):

\[\frac{36}{25}=r^2\] \[r=\frac{6}{5}\]

myininaya (myininaya):

So now you can find the area

OpenStudy (anonymous):

yup

myininaya (myininaya):

@A.Avinash_Goutham how did you get is so fast? do you know a shorter way?

OpenStudy (lgbasallote):

lol wow that sounds...uhmm long @_@

OpenStudy (anonymous):

hmmm u just need the radius....it's the length of projection of that point on to that line

OpenStudy (lgbasallote):

the radius is perpendicular to the tangent line :p

OpenStudy (anonymous):

u got a formula for that .... |ax+by+c|/(a^2 + b^2)^0.5

myininaya (myininaya):

Calculus is sweet though :)

OpenStudy (anonymous):

nd thanks for the medal :)

OpenStudy (lgbasallote):

actually it's just 4x - 3y = 5 4x - 5 = 3y slope is 4/3 that meas the slope of the radius is -3/4 since the center was given.... point slope form: y - 3 = -3/4 (x - 2) 4y - 12 = -3x + 6 4y + 3x = 18 find the intersection (3x + 4y = 18) 4 (4x - 3y = 5) -3 12x + 16y = 72 -12x + 9y = -15 im getting weird things now o.O

OpenStudy (lgbasallote):

that's where i get confused :/

OpenStudy (lgbasallote):

im trying to get the intersection of the radius and the tangent line so i can find the endpoint of the radius but it gets screwed up :C where did i go wrong @myininaya ?

myininaya (myininaya):

you are doing good actually

myininaya (myininaya):

almost there

myininaya (myininaya):

Find (x,y) then plug em into \[(x-2)^2+(y-3)^2=r^2\]

OpenStudy (lgbasallote):

yeah that's the problem cant find (x,y) :/ i get big numbers

OpenStudy (anonymous):

y = mx + c nd m =-3/4 nd u have that point

myininaya (myininaya):

\[25y=57\] \[y=\frac{57}{25} \]

myininaya (myininaya):

@lgbasallote did you get this for y?

OpenStudy (lgbasallote):

yeah

myininaya (myininaya):

\[3x+4(\frac{57}{25})=18\]

myininaya (myininaya):

I plug this into one of the equations you gave

myininaya (myininaya):

now we solve for x

myininaya (myininaya):

\[3x+\frac{228}{25}=18\] \[3x=18-\frac{228}{25}=\frac{222}{25}\] \[x=\frac{1}{3} \cdot \frac{222}{25}=\frac{74}{25}\]

myininaya (myininaya):

\[(\frac{74}{25}-2)^2+(\frac{57}{25}-3)^2=r^2\]

OpenStudy (lgbasallote):

lol i just noticed the equation of the circle is also distance formula

myininaya (myininaya):

\[\frac{576}{625}+\frac{324}{625}=r^2\] \[\frac{900}{625}=r^2\] \[\frac{36}{25}=r^2\] lol @lgbasallote

OpenStudy (lgbasallote):

6/5...oh wow they're the same *_*

OpenStudy (lgbasallote):

i thought i got offcourse when i got big numbers haah

myininaya (myininaya):

well there are more than one point on a circle lol

OpenStudy (lgbasallote):

but i used the intersection of the tangent and the radius...

OpenStudy (experimentx):

To add one more method: 4x - 3y = 5 => y = (4x-5)/3 (x - 2)^2 + (y-3)^2 = r^2 <--- put the value of y (x - 2)^2 + ((4x-5)/3-3)^2 = r^2 Now you have quadratic equation in x ... if the line was tangent ... then there must be one and only one x ... and for that ... discriminant must be zero ... i.e. b^2 - 4ac = 0 <---- there are no axes here .. only r, solve for r. If you are doing analytical geometry .. you need this trick often!!

OpenStudy (lgbasallote):

nahh im not doing analytic geometry :P you've seen my calculus problems...im just reviewing these for when i need them in calc...i suck at graphs =_=

myininaya (myininaya):

seriously it sucks i can't give more than one medal :(

OpenStudy (experimentx):

never mind ... just trade of tricks!!